Evening Shift 11/24/99

Arun Tripathi (arun@physics.ucla.edu)
Wed, 24 Nov 1999 22:35:02 -0800 (PST)


Shift Summary for eve shift, 24 Nov 99.

On Shift: Wah, Arun

Hours of beam available at all intensities: 45 min
Hours of beam available at nominal intensity: ~20 min
Hours of E799 data (special runs not included): 15 min
Hours of special runs (laser scans, muon runs...): ~8 min

Accounting of lost time: stoping the mu-run, Resetting Kumquats and
Fresh starting.

========================================================================
Runs ended this shift:
Run 15050 Muon Run

Tapes brought to FCC this shift:
(including tapes for online split)

none
========================================================================
Accesses during shift:

A few. Wah and Arun make an access to swap a preamp board
in DC4Y, Julie, Sydney and Jason go on access to fix channel 920 (no
dynode signal). Interlock drops inexplicably during Wah-Arun access.

========================================================================
Overview of shift:
No beam until 11:15. Started with muon run. Had some problems initing
a BEAM_799 run, but we are running smoothely now.

> ========================================================================
Description of Shift:



run spill time comment
------------------------------------------------------------------------
15:30 No beam.
17:20 Notice that DC4Y channel 74 is dead in the
cosmics illumination plot. Wah and Arun
make an access to fix the problem. Preamp card
was swapped, and then we discovered that
all channels corresponding to that board were
dead. Went downstairs again. Interlock drops
for no good reason, and Julie feels vindicated.
The new problem with DC4 was a blown fuse. After
replacing the preamp fuse, DC4Y looks just fine.

CsI chhannel 2446 has ped drift of -4.6 counts.
Julie updates the CSQ6 and CSPD constants, and
updated gain drift ref. plot.

19:10 We are searched and secured and interlocked again.

23:00 CsI channel 920 has no dynode signal. Julie,
Sydney and Jason make an access to look into
it. Reseating the cable at ET fixes the
problems.

Julie failed to drop the interlock this time, and
MCR people promptly made a not of it in their
e-log.
23:15 Beam comes back. Started with muon run 15050
23:20 Stopped muon run. DID NOT update cpeds, although
the run window asks the question.

23:45 Problems with initializing a BEAM_799 run
due to hung kumquats. Reset them by hand, and
did a fresh start. BEAM_799 run 15052 started.

24:00 Updated the CsI ref plots. Should take effect
from the next run.

========================================================================
Follow up Notes on Previous Shift


========================================================================